Given that HCF (306, 657) = 9, find LCM (306, 657).

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

HCF ×LCM = Product of the two given numbers
Therefore
9×LCM = 306×657
LCM = (306×657)/9 = 22338
Hence, LCM(306,657) = 22338

​Thanks
Ram Kushwah
3 years ago
Let a=306 and b=657
 
given HCF(306,657) = 9
We know that LCM x HCF = a x b
So LCM = (a x b) / HCF
=306 x 657 /9
=201042 / 9
=22338
 
Thus LCM( 306,657) =22338
 
Thanks 
Please aprove